How large should the charge of each sphere be so that the repulsive electric force between the attractive gravitay force between them. Mass of sphere= 8kg
1
Expert's answer
2021-08-27T14:39:29-0400
"k\frac{q^2}{r^2}=G\frac{m^2}{r^2}\\q=m\sqrt{\frac{G}{k}}
\\q=8\sqrt{\frac{6.67\cdot10^{-11}}{8.99\cdot10^{9}}}=6.89\cdot10^{-10}\ C"